# Lightning declines over shipping lanes following regulation of fuel sulfur emissions

### v1

@@ Abstract
By acting as cloud condensation nuclei (CCN), aerosol particles influence clouds and, in turn, the Earth’s energy balance. These aerosol interactions with clouds represent a significant uncertainty in our understanding of the Earth’s climate (1). Shipping lanes can be used as a natural experiment to reduce that uncertainty. Fuel combustion by maritime shipping vessels in the open ocean leads to the emission of aerosol particles and associated precursors, such as SO2, into relatively clean marine air. These emissions can perturb low-level marine stratus cloud droplet number distributions and related macrophysical properties, such as cloud albedo and lifetime, by increasing CCN concentrations.(2, 3, 4, 5).
Deep convective cloud (DCC) systems, commonly known as cumulonimbus or thunderstorms when lightning is generated, occur throughout the tropics, and are essential to the Earth’s water and energy cycles. In the tropics and some mid-latitude regions, the majority of precipitation and extreme weather is associated with DCCs (6). However, there is no strong consensus on the mechanisms or magnitudes of aerosol particle impacts on DCCs (7, seinfeld_improving_2016, 8, 9). Thornton et al (10) documented a potential case of persistent maritime aerosol-DCC interactions analogous to the stratocumulus ship tracks, with the discovery of enhancements in lightning located over the two busiest shipping lanes in the world, which pass through the Indian Ocean and South China Sea (Fig. 1). Lightning results from cloud electrification, which in turn requires sufficient updraft velocity, vertical ice fluxes, and super-cooled liquid water in the mixed-phase region of a DCC to induce and sustain charge separation (11, 12). Several relevant mechanisms have been proposed by which aerosol particle could enhance lightning frequency over the polluted shipping corridors, all of which involve additional aerosol particles from ship emissions causing enhanced cloud droplet nucleation (13), which subsequently leads to either 1) a perturbation to the super-cooled liquid water and ice hydrometeor distributions in the mixed-phase region of DCC (14, 15, 16); or 2) an increase in the frequency or intensity of deep convection due to changes in the vertical distribution of humidity (17) or heating (18, 19). Some combination of 1 or 2 is also possible.
In January 2020, the International Maritime Organization (IMO) reduced the amount of allowable sulfur in fuel by a factor of seven, from 3.5% to 0.5% to curb effects of the maritime shipping industry on air pollution (20). After similar regulations of fuel sulfur content, a study of ship plumes in the Baltic Sea documented a shift in the aerosol size distribution to smaller sizes (21). Recent analyses of shallow stratocumulus marine clouds over shipping lanes find changes to cloud brightness, droplet number, and droplet size, amounting to a globally averaged radiative forcing perturbation of O(0.1 Wm-2) associated with the IMO regulation, presumably due to the shift in aerosol size distribution (22, 3, 23). To date, no study has quantified the impact of the IMO regulation on DCCs.
Results
Lightning stroke density, measured by the World Wide Lightning Detection Network (WWLLN) over both the Indian Ocean and South China Sea shipping lanes (hereafter ”the shipping lanes”), has decreased since the onset of the 2020 IMO regulation of sulfur emissions (Figure 1). Relatively constant ship traffic (Figure S1) (24), paired with continuous observations of lightning from WWLLN, provide large samples to generate robust statistics. The Port of Singapore accounts for roughly 20% of the world’s bunkering fuel demand, and the two primary shipping lanes it services have nearly an order of magnitude higher traffic than other shipping lanes in the region (and around the world) (top panel) (25). We show that there is an enhancement in mean lightning stroke density associated with these shipping lanes, consistent with previous work of Thornton et al, but updated for the time period of 2010 to 2019 in Figure 1 (middle panel). Since 2020, lightning over the shipping lanes has declined by about 1 stroke km-2 year-1 as indicated by the difference in lightning stroke density between the mean of years 2020-2023 and that of 2010-2019 shown in Figure 1 (bottom panel). While some of the largest absolute declines in lightning since 2020 occur over the shipping lanes, lightning has increased or decreased in other parts of this region, suggesting that changes in the dynamic and thermodynamic context for convection over these shipping lanes should be taken into account.
Figure 1: (top) Map showing the total number of Automatic Indentification Systems (AIS) transponder calls from 2015-2021, used by maritime vessels for collision avoidance. Data from the IMF World Seaborne Trade dataset (26). (middle) Climatological mean lightning stroke density near the Port of Singapore (2010-2019). (bottom) Difference of the post-regulation period (2020-2023) lighting stroke density from the 2010-2019 climatology above
We first examine changes in the lightning stroke density enhancement over the shipping lanes using two basic controls on the background meteorology: 1) we only consider samples of precipitating clouds, using IR and radar-calibrated microwave precipitation measurements from integrated multi-satellite retrievals (IMERG) (27, 28, 29); and 2) we restrict analyses to seasons for which the conditions for lighting are favorable in each region (November to April in the Indian Ocean; June to November in the South China Sea). Given these controls, we then composite the observations as a function of distance to the shipping lanes, the center of which we define as the peak in shipping emissions from the EDGAR emissions inventory (see Methods). Mean lightning as a function of distance from the shipping lanes before (pre-IMO) and after (post-IMO) the 2020 regulation shows that the clear enhancement, between approximately 150km south to 150km north of the shipping lanes, has decreased since the regulation (Figure 2a). As before, lightning has also fallen region-wide, particularly to the south of the shipping lanes, suggesting changes to the background (non-aerosol) drivers of lightning.
To account for natural variability in the frequency and intensity of convection in the region, we construct an estimate of the natural (unperturbed) annual lightning by regressing the observed annual lightning at a given distance from the shipping lane against three variables known to relate to lightning stroke density: Convective Available Potential Energy (CAPE, discussed further below) and precipitation rates (30), as well as the annual mean Oceanic Niño Index (ONI). All variables are subjected to the same two controls listed above — precipitating (except ONI) and restricted to the high lightning season. The influence of El Niño is assumed to be uniform across the domain for a given year. The regressed variables explain roughly 33% of variance of the annual means. The regressed predictions for lightning are then removed from the observed annual mean, leaving the anomalous mean lightning stroke density, the time series of which is shown as a function of distance from the shipping lane in Figure 2b. Also shown in the right panel of Figure 2b is the monthly fraction of fuel sales at the Port of Singapore that are high- and low-sulfur.
Prior to the IMO regulation, essentially 100% of fuel sold at the port was the high sulfur type; correspondingly, the lightning anomaly over the shipping lane was 4.0 strokes km-2 year-1 on average and never fell below 3 strokes km-2 year-1 for more than one year at a time. Then, as evidenced by the bunkering fuel sales, adoption of the new regulation was immediate in 2020, such that high sulfur fuel has comprised less than 35% of fuel sold at the Port of Singapore. Since the adoption of lower-sulfur fuels in January 2020, the shipping lane lightning enhancements have declined significantly: the anomaly has fallen to 1.8 strokes km-2 year-1 on average, with none of the 4 years reaching higher than 2.2 strokes km-2 year-1. That is, the enhancement has declined by approximately 50%.
The preceding analysis relies on an assumption that the annual mean CAPE, precipitation, and ONI account for year-to-year variations in the regional drivers of convection and lightning. To further control for higher-frequency variations in the large-scale thermodynamic conditions, we examine the lightning enhancement in a 2-dimensional CAPE and precipitation space, populated by 3-hourly coincident observations of CAPE, precipitation, and lightning. Romps et al. 30 showed that CAPE $\times$ precipitation is a reasonable proxy for lightning frequency over land, and Cheng et al (31) showed that, after adjusting for a CAPE threshold, CAPE $\times$ precipitation is a reasonable proxy for tropical oceanic lightning frequency. CAPE provides an estimate of the energetic potential for deep convection and the associated updraft strength while precipitation rate indicates both the presence of a storm and a measure of its intensity. By partitioning 3-hourly lightning observations into CAPE and precipitation bins, we control for regional and temporal variations in the environmental conditions which influence lightning to further isolate the impact of the shipping lane on stroke density.
We compute lightning frequency in each CAPE-Precip bin using data from a region centered over each shipping lane and from reference regions adjacent to the shipping lanes, similar to those in (10) (Figure S2). We then compute a relative enhancement in lightning over the shipping lanes before and after the IMO regulation onset by taking the difference between corresponding CAPE-Precip bin-means in the shipping lane and reference box histograms. The shipping lane lightning enhancements in the CAPE-Precip space are shown in Figure 3. Before the IMO regulation (Pre-IMO), a strong shipping lane lightning enhancement existed in nearly every thermodynamic setting (e.g., in each CAPE-Precip bin, Figure 3a,d) for both shipping lanes.
Since the IMO regulation (Post-IMO), both shipping lanes exhibit significantly weaker lightning enhancements across most CAPE-Precip regimes (Figure 3b,e). The bin-by-bin differences between the Pre and Post-IMO lightning enhancement histograms for each shipping lane are shown in Figure 3 (c and f). For the vast majority of CAPE-Precip bins with statistically significant differences, the lightning enhancement Post-IMO has been weaker than that Pre-IMO. On average across all CAPE-Precip conditions, the lightning enhancement has decreased by 76% and by 47% for the Indian Ocean and South China Sea shipping lanes, respectively.(Figure 3c,f).
Total fuel sales at the Port of Singapore, which we take as a proxy for ship traffic and combustion in the shipping lanes, have been either stable or increasing since the 2020 regulation. The decline in the lightning enhancement since 2020 is therefore most consistent with the change in emissions composition in 2020 associated with the IMO regulation. If decreasing sulfur emissions over the shipping lanes have reduced the number of viable CCN and disrupted the mechanism for lightning enhancement, there should be a corresponding change in warm cloud microphysics. To test for such a change, we use Moderate Resolution Imaging Spectroradiometer (MODIS) satellite observations of cloud droplet number (Nd) in warm (shallow) clouds over the Indian Ocean shipping lane (where the influence of land is weaker and ship emissions are stronger) during the high-lightning season. The retrievals of Nd follow the method outlined in Zhu et al (32), such that we sample only the brightest, most active warm clouds to ensure that assumptions of adiabatic ascent are as realistic as possible. Due to the optical thickness of DCC, these retrievals can only be done for shallow warm clouds, and therefore this analysis samples a different set of conditions than the lightning observations. We thus assume that the behavior of Nd in shallow cumulus clouds from the same region is related to, though not necessarily a direct proxy for, Nd at cloud base in DCC. Unlike WWLLN lightning observations, which are influenced by the frequency of deep convection and therefore generally trend toward the ITCZ, the Nd retrievals mostly represent perturbations to CCN below shallow cumulus, consistently trending toward land.
In Figure 4, we show that prior to the IMO regulation, Nd generally increases toward land (north), with a clear enhancement in Nd over the shipping lane. This Nd perturbation is roughly 5-10% above the background (here taken simply as the average of droplet concentrations 150km north and 150km south), which is similar in magnitude to the shipping lane perturbation detected by Diamond (2) in stratoculumus clouds in the Southeast Atlantic. Since the IMO regulation, Nd has declined region-wide, and the enhancement in Nd over the shipping lane has become essentially undetectable. This corresponding decline in Nd over the shipping lane establishes additional support for the declining lightning enhancement being related to a shift in aerosol particle number-size distributions, and consequently CCN distributions over the shipping lanes induced by the IMO regulation.
Figure 4: Warm cloud-base droplet number (Nd) concentrations over the Indian Ocean, derived from MODIS observations of optical depth and effective radius following the procedure from (32). The pre-IMO regulation period is 2010-2019. Note the general increasing trend northward through the domain toward greater land influence, and the broad localized enhancement over the shipping lane prior to the regulation. Shading represents ±2SE.
Implications
We find that a previously identified enhancement in lightning stroke density over the two major shipping routes near the Port of Singapore has declined by over 40% since 2020, when the IMO regulation aimed at reducing sulfur emissions from maritime shipping came into effect. Controlling for natural variations in environmental conditions, we observe that a decline in the lightning enhancement has taken place across a range of thermodynamic regimes. The statistically robust response of lightning coincident with the onset of the IMO regulation suggests that the mechanism for the enhancement is mostly aerosol-mediated. While the ships themselves do likely act as attractors of lightning due to their prominence over the flat ocean (33), there is not evidence of a significant change in the number of ships traversing these shipping lanes over this time period (Figure S1). Further, that there was an independently observed perturbation to Nd over the Indian Ocean shipping lane prior to the IMO regulation that has since nearly vanished is indicative of a coincident change in CCN over the region. The correlated behavior between Nd and lightning lends credence to the set of CCN-mediated hypotheses proposed for invigoration of lightning (18, 17, 14).
Precisely how elevated Nd might lead to enhanced lightning remains unresolved. Both the pre-IMO and post-IMO perturbations to Nd are smaller than the enhancements of lightning, which may suggest non-linear relationships between Nd, secondary ice processes, and charge separation, or additional interactions between DCC and aerosol acting as ice nucleating particles (INP) that covary with the CCN that enhance Nd. Moreover, we currently lack the in situ observations of aerosol size distributions needed to assess how the IMO regulations have led to changes in CCN in these regions for the conditions of DCC. Given the strong updrafts in DCC, aerosol particles can experience very high super-saturation with respect to liquid water, enabling ultrafine aerosol to act as CCN, and potentially enhancing lightning without an enhancement of Nd in the less vigorous, warm clouds detectable by our methods.
Seppala et al (21) find that, in the case of a factor of 10 reduction in sulfur in ship fuel, in-plume aerosol size distributions generally shift toward lower sizes and concentrations of ultrafine particles decrease very little, or even increase as a result of the IMO regulation. As such, if the observed shipping lane lightning ehancement were the result of ultrafine particles invigorating convection (18), we would expect little change in the lightning enhancement with the IMO regulation, which is inconsistent with the decrease in lightning stroke density that we observe. We therefore conclude that the lightning enhancement over the shipping lane prior to the IMO regulation is more likely caused mostly by fine mode aerosol perturbations to 1) cloud microphysics, such as elevated supercooled liquid water concentrations or rime splintering (14) or 2) invigoration of updrafts by heightened free tropospheric humidity in polluted clouds wrought by warm rain suppression (17). That said, enhanced ultrafine particle activation to cloud drops in DCC over the shipping lanes could be an explanation for the persistent, but lower, lightning enhancement over the shipping lanes post-IMO.
We use the regulation of ship fuel sulfur to identify and assess connections between maritime shipping emissions, aerosol particle size and composition, DCC microphysics, and lightning stroke density. Further work is needed to clarify these connections and to quantify the relative roles of dynamic and microphysical responses, but our findings suggest these regions remain a useful testbed for understanding aerosol impacts on deep convective cloud and lightning. Although satellite-based radar observations of DCC microphysics are available, the spatial and temporal coverage is not yet sufficient to test whether there have been changes since the IMO. However, as time passes, statistical strength for the comparison between the two periods will increase, enabling a broader assessment of the DCC response to aerosol particles assuming maritime shipping emissions in the region remain in compliance with the IMO regulations.
Materials and Methods
Lightning stroke density observations come from the Worldwide Lightning Location Network (WWLLN), a ground-based lightning detection network with continuous global coverage of lightning at a resolution of 10km (34). WWLLN uses very low frequency radio impulses (3-30 kHz) that, upon emission from a lightning stroke, propagates between the Earth-ionosphere waveguide and disperses into a wave train. The phase and frequency of that wave train determine the time of group arrival at three or more measurement stations, which can be used to back out the location of the stroke. While the detection efficiency for individual events is lower than satellite-based methods, continuous observations for more than a decade offer much more statistical power over our region of interest.
We use integrated multi-satellite retrievals for GPM (IMERG) precipitation rates (27) and European Centre for Medium‐Range Weather Forecasts (ECMWF) ReAnalysis‐5th Generation (ERA5) atmospheric reanalyses (35) CAPE to compare the enhancement across various thermodynamic conditions. IMERG precipitation combines microwave and radar retrievals from TRMM and the GPM constellation. In ERA5, a value for CAPE is calculated for every departing level between the surface and 350hPa as follows:
[…]
@@ Results
where $z_{dep}$ is the departing level, $z_{top}$ is the level of neutral buoyancy, $\theta_{ep}$ is the virtual potential temperature of the parcel, and $\overline{\theta}_{esat}$ is the saturation virtual potential temperature of the environment. Once CAPE has been calculated for all levels, the most unstable layer is selected. We use CAPE1/2, which is directly proportional to wmax, the theoretical maximum vertical velocity achievable at a location given the stability of the atmosphere. This follows from the proportionality between kinetic energy and the square of velocity. Further discussion of CAPE as it relates to lightning can be found in (31) and (30).
@@
Lightning in Figure 1 is shown on 0.1º x 0.1º grid, calculated from 3-hourly lightning stroke densities. For subsequent calculations of the enhancement (Figures 2-3) all data (CAPE, precipitation, and lightning) is 3-hourly and mapped to a 0.5ºN x 0.625ºE grid to minimize collocation errors and noise, and for comparison with MERRA-2 aerosol and meteorological reanalysis fields. Smoothly varying data (CAPE) is remapped bilinearly, while non-smoothly varying data (precipitation and lightning) are remapped conservatively (see (36) and sources therein for further detail on regridding practices). To provide some basic control for thermodynamic and meteorological variability, we only consider precipitating clouds (precipitation greater than 0.1mm/hr). We use data from 2010 onward, as WWLLN detection efficiency was still increasing rapidly prior to 2010. The shipping lanes are defined as regions where the Emissions Database for Global Atmospheric Research (EDGAR) PM2.5 shipping emissions are greater than 5x10-12 kg m-2 s-1 (37). To remove influence from katabatic flows and sea-breeze driven convergence, we only consider the larger blue regions outlined in Figure S2. This notably removes the straight of Malacca, a region with both very high shipping emissions and active convection. There, surface convergence from land-based precipitation outflows on Sumatra and Malaysia and the adjacent landmasses make it challenging to establish a counterfactual, given the well-known land-ocean contrast in lightning stroke rates.
@@
We utilized the ”brightest 10%” method (32, 38) to obtain reliable Nd (cloud droplet number concentration) retrievals from MODIS Aqua across our target domain from 2010 to 2023. This method involves selecting the brightest 10% of clouds within each scene to calculate Nd values for every 0.5° x 0.5° grid box. The validity of this retrieval method has been corroborated through comparisons with ship-based observations (39, 40). Nd is computed using the cloud effective radius (re) and cloud optical depth ($\tau$), as described by the equation:
@@
where k represents the volume radius ratio of cloud droplets (rv) to re (k = (rv/re)³ = 0.8). The term fad denotes the adiabatic fraction, for which we assumed a constant value of 1 in our study, due to the absence of more refined alternatives (41, 42). Cw signifies the adiabatic cloud water condensation rate within an ascending cloud parcel, expressed in grams per cubic meter per meter (g m-3 m-1). The extinction efficiency factor, Qext, is assumed to be 2, and $\rho$w is the density of water. To enhance the accuracy of our Nd estimations for each 0.5° x 0.5° grid box, we excluded pixels where the solar zenith angle exceeded 65 degrees (Grosvenor and Wood, 2014). We also excluded of scenes containing mixed-phase, ice, or multilayer clouds. Consequently, after applying these filtering criteria, the remaining dataset comprised less than 1% of multilayer cloud pixels in any given grid. We use only the Indian Ocean shipping lane to maximize signal-to-noise, as the South China Sea has a much weaker signal due to its proximity to land and lower ship emissions. Inclusion of the South China Sea in the analysis does not alter the results.
Data, Materials, and Software
ERA5 CAPE may be downloaded using the Copernicus API at cds.climate.copernicus.eu. IMERG Precipitation is available for download at disc.gsfc.nasa.gov. WWLLN data is available at WWLLN.net. ONI index is available at psl.noaa.gov/data/correlation/oni.data. MODIS Aqua (MYD06) retrievals are available at ladsweb.modaps.eosdis.nasa.gov. Global ship traffic density is available at: datacatalog.worldbank.org/search/dataset/0037580/Global-Shipping-Traffic-Density. Analysis and plotting available at 10.5281/zenodo.11373991 (43)
@@ Acknowledgements
Supplemental Information for: Lightning declines over shipping lanes following regulation of fuel sulfur emissions
Figure S1: Total fuel sales at the Port of Singapore have been generally increasing in this region
Figure S2: Climatology as in Figure 1, but with the boxes used for CAPE-Precipitation enhancement calculations (black). Our approach differs slightly from that of Thornton et al., (2017), in that only the Southern Indian Ocean reference region is used for CAPE–Precip analysis due to its similarity to the Indian Ocean shipping lane in frequency of convection and precipitation during the high-lightning season (November to April). Conducting the same analysis with the Northern Indian Ocean reference region does not alter the conclusions of this analysis. Blue boxes show regions used for compositing lightning and Nd as a function of distance from the shipping lane
@@
Figure S4: Frequency of cold echoes (25dBZ or greater, above 8km), discussed in Blossey et al (2018) and citations therein as a proxy for lightning. The peak in reflectivity has diminished since the regulation, but weak sampling power does not allow us to confirm this change.

### v2

@@ Introduction
where $z_{dep}$ is the departing level, $z_{top}$ is the level of neutral buoyancy, $\theta_{ep}$ is the virtual potential temperature of the parcel, and $\overline{\theta}_{esat}$ is the saturation virtual potential temperature of the environment. Once CAPE has been calculated for all levels, the most unstable layer is selected. We use CAPE1/2, which is directly proportional to wmax, the theoretical maximum vertical velocity achievable at a location given the stability of the atmosphere. This follows from the proportionality between kinetic energy and the square of velocity. Further discussion of CAPE as it relates to lightning can be found in (Cheng et al.,, 2021) and (Romps et al.,, 2018).
@@
Lightning in Figure 1 is shown on 0.1º x 0.1º grid, calculated from 3-hourly lightning stroke densities. For subsequent calculations of the enhancement (Figures 2-3) all data (CAPE, precipitation, and lightning) is 3-hourly and mapped to a 0.5ºN x 0.625ºE grid to minimize collocation errors and noise, and for comparison with MERRA-2 aerosol and meteorological reanalysis fields. Smoothly varying data (CAPE) is remapped bilinearly, while non-smoothly varying data (precipitation and lightning) are remapped conservatively (see Staff, (2014) and sources therein for further detail on regridding practices). To provide some basic control for thermodynamic and meteorological variability, we only consider precipitating clouds (precipitation greater than 0.1mm/hr) during the high-lightning season (see SI).
We use data from 2010 onward, as WWLLN detection efficiency was still increasing rapidly prior to 2010. The shipping lanes are defined as regions where the Emissions Database for Global Atmospheric Research (EDGAR) PM2.5 shipping emissions are greater than 5x10-12 kg m-2 s-1 (Crippa et al.,, 2016). To remove influence from katabatic flows and sea-breeze driven convergence, we only consider the larger blue regions outlined in Figure S2. This notably removes the straight of Malacca, a region with both very high shipping emissions and active convection. There, surface convergence from land-based precipitation outflows on Sumatra and Malaysia and the adjacent landmasses make it challenging to establish a counterfactual, given the well-known land-ocean contrast in lightning stroke rates Cheng et al., (2021); Romps et al., (2018).
@@
We utilized the ”brightest 10%” method (Zhu et al.,, 2018; Cao et al.,, 2023) to obtain reliable Nd cloud droplet number concentration retrievals from MODIS Aqua across our target domain from 2010 to 2023. This method involves selecting the brightest 10% of clouds within each scene to calculate Nd values for every 0.5° x 0.5° grid box. The validity of this retrieval method has been corroborated through comparisons with ship-based observations (Efraim et al.,, 2020; Wang et al.,, 2021). Nd is computed using the cloud effective radius (re) and cloud optical depth ($\tau$), as described by the equation:
@@
where k represents the volume radius ratio of cloud droplets (rv) to re (k = (rv/re)³ = 0.8). The term fad denotes the adiabatic fraction, for which we assumed a constant value of 1 in our study, due to the absence of more refined alternatives (Bennartz and Rausch,, 2017; Grosvenor et al.,, 2018). Cw signifies the adiabatic cloud water condensation rate within an ascending cloud parcel, expressed in grams per cubic meter per meter (g m-3 m-1). The extinction efficiency factor, Qext, is assumed to be 2, and $\rho$w is the density of water. To enhance the accuracy of our Nd estimations for each 0.5° x 0.5° grid box, we excluded pixels where the solar zenith angle exceeded 65 degrees (Grosvenor and Wood, 2014). We also excluded of scenes containing mixed-phase, ice, or multilayer clouds. Consequently, after applying these filtering criteria, the remaining dataset comprised less than 1% of multilayer cloud pixels in any given grid. We use only the Indian Ocean shipping lane to maximize signal-to-noise, as the South China Sea has a much weaker signal due to its proximity to land and lower ship emissions. Inclusion of the South China Sea in the analysis does not alter the results. Finally, in order to remove the impact of dust storms advected over the Bay of Bengal, and to thereby reduce interannual variability in Nd outside the shipping lanes, we collocate 3-hourly MERRA-2 aerosol reanalysis output of dust and black carbon with the MODIS Nd retrievals. We then remove any Nd retrievals where dust concentrations increase above 1 ng m33{-3}- 3 or black carbon concentrations above 0.1 ng m33{-3}- 3 (approximately the 50th percentile in each case).
Data availability
ERA5 CAPE may be downloaded using the Copernicus API at cds.climate.copernicus.eu. IMERG Precipitation and MERRA-2 aerosol are available for download at disc.gsfc.nasa.gov. WWLLN data is available at WWLLN.net. ONI index is available at psl.noaa.gov/data/correlation/oni.data. Precipitation Feature reflectivity datasets are available for download at: https://atmos.tamucc.edu/trmm/data/. MODIS Aqua (MYD06) retrievals are available at ladsweb.modaps.eosdis.nasa.gov. Precipitation Feature reflectivity datasets are available for download at: https://atmos.tamucc.edu/trmm/data/. Global ship traffic density is available at: datacatalog.worldbank.org/search/dataset/0037580/Global-Shipping-Traffic-Density. Analysis and plotting available at 10.5281/zenodo.11373991 (Wright,, 2024).
@@ Acknowledgements
Supplemental Information
As noted in the main manuscript, fuel sales statistics from the Port of Singapore are a reasonable proxy for the amount of large ship activity in the two shipping lanes considered here. While not accounting for vessels that might fuel elsewhere prior to transiting one of the shipping lanes, it is likely that if a ship transits the Indian Ocean shipping lane to Singapore it will fuel there for a subsequent journey through the South China Sea (or a return to the Indian Ocean) and vise versa for a ship first transiting the South China Sea. Thus, that gross fuel sales for large vessels at the Port of Singapore has remained largely constant or even increased since 2020 (see Figure S1) suggests fairly minimal changes in ship traffic through these two shipping lanes. There is nothing like a 50% drop in shipping that might explain the change in lightning.
Figure S1: Total fuel sales at the Port of Singapore have been generally increasing
Our approach to lightning stroke density analysis differs slightly from that of Thornton et al., (2017), in that only the Southern Indian Ocean reference region is used for CAPE–Precip analysis due to its similarity to the Indian Ocean shipping lane in frequency of convection and precipitation during the high-lightning season (November to April). Conducting the same analysis with the Northern Indian Ocean reference region does not alter the conclusions of this analysis. Blue boxes in Figure S2 show regions used for compositing lightning and Nd as a function of distance from the shipping lane.
Figure S2: Climatology as in Figure 1, but with the boxes indicated in black that are used for calculating the shipping lane lightning enhancements as a function of CAPE and Precipitation presented in Figure 3.
Over both shipping lane regions, seasonal to subseasonal variability is dominated by the migration of the ITCZ, also called the monsoon, as well as the Madden-Julian Oscillation (MJO). The high lightning season in each region occurs when the ITCZ migrates over the shipping lane before reaching a stable location a few hundred kilometers away. This passage of the ITCZ, on top of periodic MJO-related convection, cause considerable variability in the background within a single high lightning season and across years. For this reason, we analyze the data first on an annual basis, taking an entire high lightning (monsoon) season (Figure 2 in the main manuscript), to complement the more rigorous high resolution (3-hourly) analysis shown in Figure 3 in the main manuscript.
For calculating the anomalous lightning enhancement over the shipping lanes shown in Figure 2B of the main manuscript, we regress out of the lightning stroke density the interannual variations in CAPE, Precipitation, and large scale climate variability represented by the ENSO ONI index. The predicted lightning stroke density using these variables outside of the shipping lane is shown in Figure S3 below.
In both Figure 2B and Figure 3 in the main manuscript, we remove the effects of CAPE and precipitation, which are considered reliable measures of MJO and monsoon activity and intensity (Zhang et al, 2022). These analyses also control for interannual variability of ITCZ strength as well as variations in MJO phase. Moreover, the intraseasonal variability associated with the MJO, shown in Figure S4, has no clear trend since 2020 and thus its specific role in the changes in the lightning enhancement is negligible.
SST-driven fluxes on heat and moisture may also be important for setting the stage for convection to occur and for driving instability; however, the connection between SSTs and lightning is less obvious than that for the more direct measures of instability (CAPE and precipitation) used here. For a more in-depth analysis of the SST-pattern drivers of convection and ITCZ migration in this region, we refer the reader to Zhang et al zhang_how_2022.
@@
We use the GPM precipitation feature database from Chuntao Liu’s group to probe the shipping lane enhancement of cold, strong reflectivity echoes liu_cloud_2008. Note that this approach to examining reflectively is slightly different from that done in Thornton et al. Thornton et al., (2017), but follows that in Blossey et al Blossey et al., (2018). Due to infrequent sampling and small swath width, the reflectivity signal shown here requires very long records to show the effect of the shipping lane (previous studies, such as Blossey et al Blossey et al., (2018), use the full TRMM record, which is well over a decade). Splitting the relatively short (8 year) GPM record of reflectivity into a pre- and post-IMO period therefore greatly reduces the statistical power of the retrievals. Indeed, the peak in reflectivity established over the shipping lane between 2016 and 2019 has diminished (Figure S5). However, the difference between the two periods is not statistically significant (p less than 0.05), and we therefore refrain from concluding that there has or has not been a change to reflectivity above the shipping lanes.
Figure S5: Frequency of cold echoes (25dBZ or greater, above 8km) from GPM, discussed in Blossey et al (2018) and citations therein as a proxy for lightning. The peak in reflectivity has diminished since the regulation, but weak sampling power does not allow us to confirm this change (the pre-IMO record is restricted GPM’s record, i.e. years since 2016).
While a correspondence between aerosol optical depth (AOD) and CCN has been documented in some highly polluted regions, this relationship is not established for the typically clean (AOD<0.02) regions of interest here. Therefore, as noted in the main manuscript, we expect that perturbations to CCN in the shipping lane regions is best represented by MODIS Nd retrievals, rather than by AOD because 1) clear sky conditions are rare in this region leading to limited AOD retrievals and consequently limited statistical power, 2) the background AOD is very low, such that an expected perturbation is barely above the limit of detection, and 3) AOD being proportional to aerosol cross sectional area columns does not directly correspond to CCN number at cloud base, the subject of this paper. For example, as shown in Figure S5, AOD estimates from MERRA-2 reanalysis, nudged by both their aerosol model and the limited AOD observations by MODIS, exhibit no observable enhancement over the shipping lanes prior to or after the 2020 IMO regulation. This lack of AOD enhancement is somewhat expected given that ship emissions of aerosol particles are concentrated in smaller, less radiatively active particle sizes [see, e.g., Hobbs et al., Hobbs et al., (2000) and Seppala et al. Seppälä et al., (2021)], statistical sampling is limited by infrequent cloud-free scenes, and that particle scavenging and dilution rates in these regions are relatively high due frequent convection.
Figure S6: Mean Aerosol Optical Depth from MERRA-2 as a function of distance from the shipping lane. Shading indicates 95% confidence.